# Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$

### v1

@@
We present a new determination of the star-forming main sequence (MS), obtained through stacking 100k $K$-band-selected galaxies in the far-IR Herschel and James Clerk Maxwell Telescope (JCMT) imaging. By fitting the dust emission curve to the stacked far-IR photometry, we derive the IR luminosities ($L_{\rm IR}$) and, hence, star formation rates (SFR) out to $z\sim 7$. The functional form of the MS is found, with the linear SFR-$M_{\ast}$ relation that flattens at high stellar masses and the normalization that increases exponentially with redshift. We derive the corresponding redshift evolution of the specific star formation rate (sSFR) and compare our findings with the recent literature. We find our MS to be exhibiting slightly lower normalization at $z\lesssim 2$ and to flatten at larger stellar masses at high redshifts. By deriving the relationship between the peak dust temperature ($T_{\rm d}$) and redshift, where $T_{\rm d}$ increases linearly from $\sim 20$ K at $z=0.5$ to $\sim 50$ K at $z=6$, we conclude that the apparent inconsistencies in the shapes of the MS are most likely caused by the different dust temperatures assumed when deriving SFRs in the absence of far-IR data. Finally, we investigate the derived shape of the star-forming MS by simulating the time evolution of the observed galaxy stellar mass function (GSMF). While the simulated GSMF is in good agreement with the observed one, some inconsistencies persist. In particular, we find the simulated GSMF to be somewhat overpredicting the number density of low-mass galaxies at $z\gtrsim 2$.
@@
Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$
@@
M. P. Koprowski Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$    J. V. Wijesekera Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$    J. S. Dunlop Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$    D. J. McLeod Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$    M. J. Michałowski Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$    K. Lisiecki Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$    R. J. McLure Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$
@@ 1 Introduction
The increasing amount of multi-wavelength data, obtained as part of wide-field and deep surveys, enables a progressively detailed examination of large samples of galaxies. Substantial advancements in determining important physical properties of galaxies (e.g. photometric redshifts, stellar masses, dust attenuation curves) have enabled the identification (in the rest-frame color-color space) of two separate types of galaxy populations: star-forming and quiescent (SFGs and QGs; e.g. Labbé et al. 2005; Wuyts et al. 2007; Ilbert et al. 2010). The main difference between these two distinct populations of galaxies is their star-formation activity, where it has been shown that the SFGs follow a tight correlation between their star formation rate, SFR, and stellar mass, $M_{\ast}$, out to the highest redshifts probed – the so-called star-forming main sequence (MS; e.g. Speagle et al. 2014; Tomczak et al. 2016; Daddi et al. 2022; Popesso et al. 2023). Most commonly, the SFR-$M_{\ast}$ relation can be parametrized with a simple power-law:
@@
where $\alpha$ and $\beta$ represent the slope and normalization of the main sequence, respectively. While the normalization has been shown to increase in redshift, where galaxies exhibit higher SFRs at earlier epochs, the slope of the main sequence was found to be close to unity, mostly independent of redshift, with values ranging between $\sim 0.8$ and 1.3 (e.g. Lee et al. 2015; Tomczak et al. 2016; Popesso et al. 2023). More recently, a number of studies found that the star-forming main sequence flattens at higher masses (e.g. Lee et al. 2015; Schreiber et al. 2015; Leslie et al. 2020; Leja et al. 2022; Popesso et al. 2023), reaching an asymptotic value of the SFR.
@@
The exact shape and evolution of the main sequence is still widely discussed in the literature, with many studies presenting somewhat different perspectives. The sources of the discrepancies are many, including the assumed Initial Mass Function (IMF), stellar population synthesis (SPS) modeling, star formation histories (SFHs), dust extinction curves, etc. Most importantly, different techniques used for the estimation of the total SFRs can affect the results significantly. Since part of the UV emission coming from young stars gets absorbed by the interstellar dust, it is crucial to be able to properly quantify this effect. Measuring the dust attenuation from the rest-frame UV/optical data alone produces very uncertain results, because of the degeneracy between the age (different SFHs) and reddening (different metallicities, extinction curves). The advent of the Herschel Space Observatory (Pilbratt et al. 2010) allowed for the direct detection of the energy being re-emitted by dust in the far-IR and, therefore, more reliable determination of the total (UV+IR) SFRs. However, since Herschel far-IR data suffers from relatively low resolution, only the most IR-luminous sources can be detected at high redshifts and, hence, samples of star-forming galaxies studied with Herschel tend to be incomplete, often biasing the results towards large values of the SFRs. Speagle et al. (2014) and, more recently, Popesso et al. (2023) compiled various studies in the literature in order to convert them to the same absolute calibrations in an attempt to solve some of the issues listed above. While they carefully address all the inconsistencies caused by different modelling assumptions, the uncertainties associated with the various total SFR indicators, as well as the completeness of samples compiled, remain a concern.
[…]
@@ 2.1 Optical/near-IR catalogs
We start with the optical/near-IR catalogs of the UKIDSS Ultra Deep Survey (UDS) and COSMOS fields. The details of the data used, galaxies selection process and determination of photometric redshifts and stellar masses are presented in McLeod et al. (2021), with the shortened description given below.
@@ 2.1.1 Imaging
[…]
@@ 2.1.2 Photometric redshifts and stellar masses
We constructed our parent multi-wavelength catalogs by running Source-Extractor (Bertin & Arnouts 1996) in dual-image mode, with the K-band as the detection image. Photometry was measured in 2-arcsec diameter apertures on the PSF-homogenized UV to near-IR images. For the lower resolution Spitzer IRAC 3.6 and 4.5 ${\rm\mu m}$ data, the photometry was measured using the software tphot (Merlin et al. 2015).
For each object in the catalog, the photometric redshift we assigned was the median of six SED fitting runs, employing three different software codes. With LePhare, we included runs with the Bruzual & Charlot (2003), Pegase (Fioc & Rocca-Volmerange 1999) and COSMOS (Ilbert et al. 2009) template libraries. We also included two eazy (Brammer et al. 2008) runs, using the PCA and Pegase templates, and one further run with the BPZ code (Benítez 2000), using the (Coleman et al. 1980) templates.
Stellar masses were measured for each object by fixing the photometric redshift to $z_{\rm med}$, and re-fitting the SED using LePhare with the Bruzual & Charlot (2003) library. This template set included a Chabrier (2003) initial mass function, a Calzetti et al. (2000) dust attenuation law and IGM absorption as in Madau (1995). The resulting catalogs were cleaned of stars, AGN, artefacts and objects with contaminated photometry. Finally, a further $\chi^{2}<40$ goodness-of-fit cut was implemented on the catalogs.
2.2 Far-IR
@@ 3.1 Sample selection
Since we use the derived SFR-$M_{\ast}$ relations to ‘grow’ the McLeod et al. (2021) galaxy stellar mass function in Section 5.4, we decided to follow them in choosing the mass and redshift bins, while at the same time extending our analysis to $z=6.75$. At $z>3.75$, in order to be able to reach the detection threshold, the mass and redshift bins have been accordingly stretched. The procedure for estimating the mass completeness in the UDS and COSMOS fields is explained in McLeod et al. (2021). In brief, the mass completeness was estimated using simulations, where $K$-band imaging was produced by injecting a set of artificial galaxies with a wide range of physical properties. Adopting the original set-up, the corresponding photometry catalogs were then constructed using sextractor and the 90% mass-completeness limit, defined as the stellar mass below which 90% of the limiting stellar masses lie, at each redshift was derived. The resulting redshift and stellar mass bins adopted in this work are listed in Table 2.
In order to derive the SFR-$M_{\ast}$ relation for star-forming sample, we first need to separate it from the quiescent population. Similarly to McLeod et al. (2021) we identify passive galaxies using the UVJ color-color selection (Williams et al. 2009) from Carnall et al. (2020):
@@
where $U$,$V$ and $J$ rest-frame magnitudes were generated from the best-fit SEDs as explained in McLeod et al. (2021). The advantage of this method is that it is able to separate sources where reddening is caused by aging from those where it is produced by dust extinction, thereby decreasing the chance of contaminating the dust-enshrouded star-forming sample with quiescent galaxies. Because at high redshifts the rest-frame $J$-band can no longer be traced by the photometry available in this work, the UVJ color-color selection was performed out to $z\sim 4$, as depicted in Figure 1.
Figure 1: The results of the $UVJ$ color-color selection results, defined in Equation 2 shown for redshift bins adopted in this work. The $UVJ$ colors (Williams et al. 2009) are used in order to distinguish between sources reddened due to the lack of star formation (quiescent galaxies – upper left section of the plots) from those where red colors are caused by dust extinction (star-forming galaxies – main locus in the plots).
3.2 Far-IR fluxes
Since current far-IR/submm surveys struggle to reach the (sub-)mJy depths, necessary for the detection of the dust emission in typical star-forming galaxies, we resort here to the well-known method of stacking (e.g. Whitaker et al. 2014; Schreiber et al. 2015; Tomczak et al. 2016; Koprowski et al. 2018). In order to determine the far-IR fluxes at a given band, we cut a square stamp around the position of each source, with the size of approximately 3 $\times$ the full width at half maximum (FWHM) of the corresponding beam. The average far-IR fluxes at each pixel are then evaluated following the inverse-variance weighting:

### v2

# Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$

@@
Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$
@@
M. P. Koprowski Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$ Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$    J. V. Wijesekera Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$ Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$    J. S. Dunlop Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$ Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$    D. J. McLeod Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$ Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$    M. J. Michałowski Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$ Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$ Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$ Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$    K. Lisiecki Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$ Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$ Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$ Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$    R. J. McLure Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$ Charting the main sequence of star-forming galaxies out to redshifts $z\lesssim 5.7$
We present a new determination of the star-forming main sequence (MS), obtained through stacking 100k $K$-band-selected galaxies in the far-infrared (FIR) Herschel and James Clerk Maxwell Telescope (JCMT) imaging. By fitting the dust emission curve to the stacked FIR photometry, we derive the IR luminosities ($L_{\rm IR}$), and hence the star formation rates (SFRs) out to $z\lesssim 5.7$. The functional form of the MS is found, with the linear SFR-$M_{\ast}$ relation that flattens at high stellar masses and the normalization that increases exponentially with redshift. We derive the corresponding redshift evolution of the specific star formation rate (sSFR) and compare our findings with the recent literature. We find our MS to be exhibiting slightly lower normalization at $z\lesssim 2$ and to flatten at somewhat larger stellar masses at high redshifts. By deriving the relationship between the peak dust temperature ($T_{\rm d}$) and redshift, where $T_{\rm d}$ increases linearly from $\sim 20$ K at $z=0.5$ to $\sim 50$ K at $z=5$, we conclude that the apparent inconsistencies in the shapes of the MS are most likely caused by the different dust temperatures assumed when deriving SFRs in the absence of FIR data. Finally, we investigate the derived shape of the star-forming MS by simulating the time evolution of the observed galaxy stellar mass function (GSMF). While the simulated GSMF is in good agreement with the observed one, some inconsistencies persist. In particular, we find the simulated GSMF to be slightly overpredicting the number density of low-mass galaxies at $z\gtrsim 2$.
@@ 1 Introduction
The increasing amount of multiwavelength data, obtained as part of wide-field and deep surveys, enables a progressively detailed examination of large samples of galaxies. Substantial advancements in determining important physical properties of galaxies (e.g., photometric redshifts, stellar masses, dust attenuation curves) have enabled the identification (in the rest-frame color-color space) of two separate types of galaxy populations: star-forming ones and quiescent ones (SFGs and QGs; e.g., Labbé et al. 2005; Wuyts et al. 2007; Ilbert et al. 2010). The main difference between these two distinct populations of galaxies is their star formation activity; it has been shown that the star formation rate (SFR) and stellar mass ($M_{\ast}$) of the SFGs follow a tight correlation out to the highest redshifts probed – the so-called star-forming main sequence (MS; e.g., Speagle et al. 2014; Tomczak et al. 2016; Daddi et al. 2022; Popesso et al. 2023). Most commonly, the SFR-$M_{\ast}$ relation can be parameterized with a simple power law:
@@
where $\alpha$ and $\beta$ represent the slope and normalization of the MS, respectively. While the normalization has been shown to increase in redshift, with galaxies exhibiting higher SFRs at earlier epochs, the slope of the MS has been found to be close to unity, mostly independent of redshift, with values ranging between $\sim 0.8$ and 1.3 (e.g., Lee et al. 2015; Tomczak et al. 2016; Popesso et al. 2023). More recently, a number of studies have found that the star-forming MS flattens at higher masses (e.g., Lee et al. 2015; Schreiber et al. 2015; Leslie et al. 2020; Leja et al. 2022; Popesso et al. 2023), reaching an asymptotic value of the SFR.
@@
The exact shape and evolution of the MS is still widely discussed in the literature, with many studies presenting somewhat different perspectives. The sources of the discrepancies are many, including the assumed initial mass function (IMF), stellar population synthesis modeling, star formation histories (SFHs), and dust extinction curves. Most importantly, different techniques used for the estimation of the total SFRs can affect the results significantly. Since part of the UV emission coming from young stars gets absorbed by the interstellar dust, it is crucial to be able to properly quantify this effect. Measuring the dust attenuation from the rest-frame UV/optical data alone produces very uncertain results, because of the degeneracy between the age (different SFHs) and reddening (different metallicities, extinction curves). The advent of the Herschel Space Observatory (Pilbratt et al. 2010) allowed for the direct detection of the energy being re-emitted by dust in the far-infrared (FIR), and therefore a more reliable determination of the total (UV+IR) SFRs. However, since Herschel FIR data suffers from a relatively low resolution, only the most IR-luminous sources can be detected at high redshifts, and hence samples of SFGs studied with Herschel tend to be incomplete, often biasing the results toward large values of the SFRs. Speagle et al. (2014) and, more recently, Popesso et al. (2023) compiled various studies in the literature in order to convert them to the same absolute calibrations in an attempt to solve some of the issues listed above. While they carefully address all the inconsistencies caused by different modeling assumptions, the uncertainties associated with the various total SFR indicators, as well as the completeness of samples compiled, remain a concern.
In this work, we make use of the sample of $\sim 100$k $K$-band-selected galaxies from the ground-based UV+optical+near-IR $\sim 2$ deg2 imaging of the UKIDSS Ultra Deep Survey (UDS) and COSMOS fields (McLeod et al. 2021). Galaxies are stacked in bins of redshift and stellar mass in the FIR Herschel and the James Clerk Maxwell Telescope (JCMT) SCUBA-2 (Holland et al. 2013) maps, in order to determine the shape and redshift evolution of the star-forming MS out to $z\lesssim 5.7$. The stacking approach, used by several authors before (e.g., Tomczak et al. 2016; Leslie et al. 2020; Mérida et al. 2023), has several advantages. The sample of galaxies stacked is mass-complete, assuring that the resulting values of the IR luminosity (and SFR) are not biased due to selection effects. In addition, thanks to the inclusion of the SCUBA-2 850 ${\rm\mu m}$ data, our stacked photometry maps the Rayleigh–Jeans portion of the dust emission spectrum, allowing for the precise determination of the dust temperatures, and thus the SFRs.
The paper is structured as follows. In Sect. 2, the data used in this work is described. We outline the stacking procedure and the star-forming sample selection process in Sect. 3. Section 4 explains the methodology used in the determination of the star-forming MS for our sample. We discuss the results of this work in Sect. 5. In particular, the shape and evolution of the star-forming MS and its relation to the so-called Schmidt-Kennicutt relation (SK relation; Schmidt 1959; Kennicutt 1998) is explained in Sect. 5.1, followed by a discussion of the variations in the peak dust temperature with IR luminosity and redshift in Sect. 5.2 and a comparison to other works in Sect. 5.3. In Sect. 5.4, we attempt to interpret some of the discrepancies between the different shapes of the MS found in the literature, by “growing” the observed galaxy stellar mass function (GSMF) in time. Finally, we summarize our conclusions in Sect. 6. Throughout the paper, we use the Chabrier (2003) stellar IMF with an assumed flat cosmology of $\Omega_{\rm m}=0.3$, $\Omega_{\Lambda}=0.7$, and ${\rm H}_{0}=70$ km s-1 Mpc-1.
@@ 2.1 Optical/near-IR catalogs
We started with the optical/near-IR catalogs of the UKIDSS Ultra Deep Survey (UDS) and COSMOS fields. The details of the data used, galaxies selection process and determination of photometric redshifts and stellar masses are presented in McLeod et al. (2021), with the shortened description given below.
@@ 2.1.1 Imaging
For the UKIDSS UDS field (Lawrence et al. 2007), we included UDS DR11 imaging (Almaini et al., in preparation) in the near-IR JHK bands, with Y-band imaging from the VISTA VIDEO DR4 release over XMM-LSS (Jarvis et al. 2013). We complemented this with UV imaging in $u^{\ast}$-band from CFHT MegaCam, and optical imaging in BVRiz from Subaru Suprime-Cam (Furusawa et al. 2008). We also employed imaging in the narrowband NB921 (Koyama et al. 2011) and the refurbished Suprime-Cam $z$-band ($z^{\prime}_{\rm new}$; Furusawa et al. 2016). To extend our wavelength coverage into the mid-infrared, we utilized Spitzer IRAC imaging in 3.6 $\mu$m and 4.5 $\mu$m, combining the SEDS (Ashby et al. 2013), S-CANDELS (Ashby et al. 2015), and SPLASH (PI Capak; see Mehta et al. 2018) programs. The effective area of the overlapping UDS imaging was 0.69 deg2, after masking for bright stars.
@@
Figure 1: Results of the $NUVrJ$ color-color selection results, defined in Eq. 2 shown for redshift bins adopted in this work. The $NUVrJ$ colors (Ilbert et al. 2013) are used in order to distinguish between sources reddened due to the lack of star formation (quiescent galaxies – upper left section of the plots) from those where red colors are caused by dust extinction (SFGs – main locus in the plots).
@@ 2.1.2 Photometric redshifts and stellar masses
As is explained in detail in McLeod et al. (2021), the parent multiwavelength catalogs were constructed by running Source-Extractor (Bertin & Arnouts 1996) in dual-image mode, with the K-band as the detection image. Photometry was measured in 2-arcsec diameter apertures on the PSF-homogenized UV to near-IR images. For the lower resolution Spitzer IRAC 3.6 and 4.5 ${\rm\mu m}$ data, the photometry was measured using the software tphot (Merlin et al. 2015).
For each object in the catalog, the photometric redshift we assigned was the median of six SED fitting runs, employing three different software codes. With LePhare, we included runs with the Bruzual & Charlot (2003), Pegase (Fioc & Rocca-Volmerange 1999) and COSMOS (Ilbert et al. 2009) template libraries. We also included two eazy (Brammer et al. 2008) runs, using the PCA and Pegase templates, and one further run with the BPZ code (Benítez 2000), using the Coleman et al. (1980) templates. The resulting values of $\sigma_{z}$, defined as $1.483\times{\rm MAD}(dz)$, where MAD is the median absolute deviation and $dz=(z_{\rm phot}-z_{\rm spec})/(1+z_{\rm spec})$, for COSMOS and UDS fields were found to be equal to 0.019 and 0.022, respectively (McLeod et al. 2021).
Stellar masses were measured for each object by fixing the photometric redshift to $z_{\rm med}$, and re-fitting the SED using LePhare with the Bruzual & Charlot (2003) library. This template set included a Chabrier (2003) IMF, a Calzetti et al. (2000) dust attenuation law and IGM absorption as in Madau (1995). The values were found to be in excellent agreement with those derived by the CANDELS team (see Table 2 of McLeod et al. 2021), with a tight 1:1 relation and a typical scatter of ±$\pm 0.05$ dex. The resulting catalogs were then cleaned of stars, AGN, artifacts and objects with contaminated photometry. Finally, a further $\chi^{2}<40$ goodness-of-fit cut was implemented on the catalogs.
2.2 Far-infrared
@@ 3.1 Sample selection
For the purpose of successful stacking (i.e., to maximize the number of sources in each stack), we designed each redshift bin to encompass $\sim 1$ billion years of Universe evolution, ranging from 1 to 9 billion years after the Big Bang (with the corresponding redshift bins of [0.45, 0.6, 0.75, 1.0, 1.25, 1.6, 2.2, 3.2, 5.7]; see Table 2). We also made sure that the sample we stacked is mass-complete. The procedure for estimating the mass completeness in the UDS and COSMOS fields is explained in McLeod et al. (2021). In brief, the completeness was estimated using simulations, in which $K$-band imaging was produced by injecting a set of artificial galaxies with a wide range of physical properties. Adopting the original setup, the corresponding photometry catalogs were then constructed using sextractor. In order to determine the 90% mass-completeness limit, the relation from Pozzetti et al. (2010) was used to derive the limiting stellar mass for each object in our sample. The 90% mass completeness limit, below which 90% of the limiting stellar masses lie, could then be derived as a function of redshift. The resulting redshift and stellar mass bins adopted in this work are listed in Table 2.
In order to derive the SFR-$M_{\ast}$ relation for the star-forming sample, we first needed to separate it from the quiescent population. Here, we adopted the selection based on the $NUVrJ$ colors (Ilbert et al. 2013):
@@
where $NUV$, $r$ and $J$ rest-frame magnitudes were generated from the best-fit SEDs. The advantage of this method is that it is able to separate sources for which reddening is caused by aging from those for which it is produced by dust extinction, thereby decreasing the chance of contaminating the dust-enshrouded star-forming sample with quiescent galaxies. Because at high redshifts the rest-frame $J$-band can no longer be traced by the photometry available in this work, the $NUVrJ$ color-color selection was performed out to $z\sim 4$, as is depicted in Fig. 1.
To avoid overestimating the SFRs at the high-mass end of the MS, we further detected and removed the starburst galaxy population from our stacking sample. To identify starbursts, we used the ALMA-detected sources in the COSMOS (Liu et al. 2019) and UDS (Stach et al. 2019) fields, with associated stellar masses and SFRs, as has been found by Liu et al. (2019) and Dudzevičiūtė et al. (2020), respectively. Following Elbaz et al. (2018), starbursts are defined as galaxies with ${\rm SFR/SFR_{MS}}>3$, where ${\rm SFR_{MS}}$ is the corresponding MS value, given the redshift and stellar mass of each ALMA-detected source, with the functional form of the star-forming MS of Popesso et al. (2023).
3.2 Far-infrared fluxes
Since current FIR/submillimeter surveys struggle to reach the (sub-)millijansky depths necessary for the detection of the dust emission in typical SFGs, we have resorted here to the well-known method of stacking (e.g., Whitaker et al. 2014; Schreiber et al. 2015; Tomczak et al. 2016; Koprowski et al. 2018). In order to determine the FIR fluxes at a given band, we cut a square stamp around the position of each source, with the size of approximately 3 $\times$ the full width at half maximum (FWHM) of the corresponding beam. The average FIR fluxes at each pixel were then evaluated following the inverse-variance weighting: